2) Determine the tension in each rope-cable if the weight of the crate is 1500 lb. z 2 ft 2 ft B 4 ft 5.5 ft 2.5 ft 6 ft
Answer #1

The solution for calculating tension in each rope cable is given below as image file. 28 B C 1467 TAS D X 6th ..0:30.6 -0.9231 +0.2316) 2) weight of trake (W) = 150015 TACK SSH А FDA Crano N 2.5ft het us find uUAL Date S M T W T F S in ) Notes m2-fo=87-0-8545°40-4297 TAL AD b=0 Ditt 6;²+2.50 DAL=6:5 TUDA of +0.9233+0.3052 Sum of forc-16AB - 2.038 (1066) AB = -3896-104 3.772 AB -38 96 104 AB 1,032.901) > Answer AC = 1066an) - 1056 x 61032 901). AC 1,101. 07
Answer #2

Z 2f1f6 -Y (1) 42 (k) B с 4 ft FABIO 5.5ft +xli) FAC 0 FAD 1500 lb 2.5ft 6 ft -X (i) * ZWk) Figure [(a)] +46) - •> Let Fas &= 6.5 TAB! .. AB > Vector AB - Z - Ā = (-2î + 0,^+4 k) - 10+6;4+2:56) AB = -21 166+1.5 Ê Magnitude of TAB = LABI = x+y*+z2 -√step2 : To find cartesian rectors i Let F, Ę, ç, be the cartesian vectors. FAB F = UAB. FAB F = (-0.308 î - 0.923ỹ +0.231 k).Finally equating the i, & ħ components to zero => [(-0.308 Fe + 0.286 FADA] = 0 .-0-308 F A 13 +0.286 E AC - 0 Also, [€0.923
Answer #3

Tension vectors;

\vec{T}_{AB} = T_{AB}\left [ \frac{\vec{r}_{AB}}{|r_{AB}|} \right ] = T_{AB}\left [ \frac{-2\hat{i}-6\hat{j}+(4-2.5)\hat{k}}{\sqrt{2^2+6^2+(4-2.5)^2}} \right ]

\Rightarrow \vec{T}_{AB} =-0.3077T_{AB}\hat{i}-0.9231T_{AB}\hat{j}+0.2308T_{AB}\hat{k}

and

\vec{T}_{AC} = T_{AC}\left [ \frac{\vec{r}_{AC}}{|r_{AC}|} \right ] = T_{AC}\left [ \frac{2\hat{i}-6\hat{j}+(5.5-2.5)\hat{k}}{\sqrt{2^2+6^2+(5.5-2.5)^2}} \right ]

\Rightarrow \vec{T}_{AC} =0.2857T_{AC}\hat{i}-0.8571T_{AC}\hat{j}+0.4286T_{AC}\hat{k}

Force vector in the strut;

\vec{F}_{DA} = F_{DA}\left [\frac{\vec{r}_{DA}}{|r_{DA}|} \right ] = F_{DA}\left [ \frac{6\hat{j}+2.5\hat{k}}{\sqrt{6^2+2.5^2}} \right ]

\Rightarrow \vec{F}_{DA} =0.9231F_{DA}\hat{j} + 0.3846F_{DA}\hat{k}

Weight;

\vec{W}=-1500\hat{k}

By equilibrium of forces;

\sum F_x = 0

\Rightarrow -0.3077T_{AB}+0.2857T_{AC}=0 \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;...(1)

and

\sum F_y = 0

\\\Rightarrow -0.9231T_{AB}-0.8571T_{AC}+0.9231F_{DA}=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;...(2)

and

\sum F_z = 0

\\\Rightarrow 0.2308T_{AB}+0.4286T_{AC}+0.3846F_{DA}-1500=0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;...(3)

Solve (1), (2) and (3);

T_{AB} = 1026\;\;lb

T_{AC} = 1105\;\;lb

...(Answer)